# Smeared phase transition in the dissipative random quantum Ashkin-Teller model

### v1

@@ I Introduction
Quenched disorder underlies a manifold of profound effects near quantum phase transitions: For instance, disorder-induced inhomogeneities lead to exotic universality classes such as those typified by the infinite randomness critical point (fisher_1; fisher_2), and the strong off-critical Griffiths singularities (griffiths_original; mccoy-prl69). Furthermore, in certain situations these inhomogeneities could lead to the destruction or smearing (smeared_hoyos_1; smeared_hoyos_2) of the putative quantum phase transition. The moorings of the effects described above is centered around the so-called rare regions (RRs): arbitrarily large regions in space which, due to a statistically rare configuration of the quenched-disordered couplings, is locally ordered and weakly coupled to the bulk. These RRs, characterized by slow dynamics, dominate the low-energy behavior of physical observables (glass_1; glass_2), (for reviews, see Refs. (vojta2006rare; igloi-review; vojta2013phases; igloi-monthus-review2)).
@@
In practice it is important to realize that real materials are seldom described as closed systems. Thus, it is of paramount importance to understand how the phases and phase transitions gets affected when the system is connected to a dissipative reservoir. The combined effects of quenched disorder and dissipation have been investigated in the Griffiths phase where it was established that sufficiently strong dissipation can completely damp the fluctuations of a RR (millis-morr-schmalian-prl01; vojta-prl03; vojta-schmalian-prb05; vojta-hoyos-prl14). Further, Strong-Disorder Renormalization-Group (SDRG) based analytical study established that for the case of the Random Transverse-Field Ising Model (RTFIM), in any dimension, the critical point is entirely destroyed by smearing when dissipation is Ohmic or Subohmic (smeared_hoyos_1; smeared_hoyos_2). Physically, the smearing phenomena occurs because dissipation freezes the dynamics of sufficiently large RRs, independent of the bulk, i.e., each rare region orders independently from the bulk, leading to local phase transitions that can take place at different values of the control parameter. Thus, the paramagnetic Griffiths phase becomes an inhomogeneously ordered ferromagnetic phase. Notably, experimental evidence of smeared phase transitions has been observed in the compounds Sr1-xCaxRuO3 (experimental_1) and CePd1-xRhx (experimental_2). More recently, smeared phase transitions have also been identified in a numerical study of strange metals (strangemetals), and have also potentially seen in low-dimensional superconductors (kaur-etal-njp24).
@@
This work belongs to the genre of problems that we have discussed in the previous paragraph: Namely, we study the dissipative random quantum Ashkin-Teller (RQAT) model. Besides the usual paramagnetic (PM) and ferromagnetic (FM) phase, this model features an additional phase, called the product phase, which exhibits a composite (or intertwined) order (1981_ashkinteller). We inquire whether the transitions between this product phase and the other phases are smeared by the combined effects of dissipation and disorder. Interestingly, we show that Ohmic dissipation smears only the transition between the product and FM phases, while the transition between product and PM phases remains sharp. This conclusion is robust and persists across different forms of system-bath couplings. Surprisingly, even though dissipation is coupled to the original order parameter fields, it does not couple generically to the intertwined order parameter field and, therefore, does not damp the quantum fluctuations of the associated rare regions that form near the transition line between the product and PM phases.
@@
Finally, besides this fundamental motivation, it is worth mentioning that the Ashkin-Teller model has applications in other contexts such as to describe layers of atoms absorbed on surfaces (bak-etal-prl85), current loops in high-Tc superconductors (aji-varma-prl07), and the elastic response of DNA molecules (chang-wang-zheng-ctp08). The understanding of dissipation in these contexts is another motivation for our work.
@@
In the remainder of this work, we review the effects of disorder on the quantum Ashkin-Teller chain in Sec. II, introduce the model studied in Sec. III, apply the adiabatic renormalization-group and the strong-disorder renormalization-group methods to our problem in Secs. IV and V, respectively. The resulting phase diagram is then discussed in Sec. VI. Finally, we give concluding remarks in Sec. VII. Other possible models of dissipation are considered in Appendix A and we find that our conclusions remain qualitatively correct.
II Review of the disspationless model
@@
The Hamiltonian of the one-dimensional random quantum Ashkin-Teller (RQAT) model is (1943_ashkinteller; 1981_ashkinteller; 2001_ashkinteller; 2008_ashkinteller; 2014_ashkinteller)
@@
where $S_{\nu,i}^{z}$ and $S_{\nu,i}^{x}$ are Pauli matrices representing local spin-1/2 moments. The first two terms on the right-hand-side of Eq. (1) describe two independent random transverse-field Ising (RTFI) chains whereas the remaining terms quantify the couplings between them. The index $\nu=1,2$, referred to as the color index, distinguishes the two chains and the index $i$ distinguishes different lattice sites. The interactions $J_{i}$ and transverse fields $h_{i}$ are strictly positive, independent random variables. The coupling strengths are characterized by the ratios $\epsilon_{h,i}=g_{i}/h_{i}$ and $\epsilon_{J,i}=K_{i}/J_{i}$. For simplicity, we assume that the (bare) values $\epsilon_{h,i}=\epsilon_{J,i}=\epsilon_{I}>0$, i.e., they are uniform and site independent. However, they flow under the renormalization-group transformations and become random variables (2014_ashkinteller). The low-energy physical behavior of the model (1) can be divided into two regimes: the weak-coupling regime ($\epsilon_{I}<1$) and the strong-coupling regime ($\epsilon_{I}>1$). In the low-energy limit of the weak-coupling regime, the two Ising chains effectively decouple (2001_ashkinteller; 2008_ashkinteller) and the resulting physics is that of two identical RTFI chains. Therefore, there is a direct quantum phase transition between a ferromagnet (where the quantum average $\left\langle S_{\nu,i}^{z}\right\rangle\neq 0$) and paramagnet (where $\left\langle S_{\nu,i}^{z}\right\rangle=0$). The transition, which occurs when $\prod J_{i}=\prod h_{i}$, is of infinite-randomness type and is surrounded by the Griffiths ferro- and paramagnetic phases. In the strong-coupling regime $\epsilon_{I}>1$, on the other hand, this transition is through an intervening product phase exhibiting a composite order, i.e., $\left\langle S_{\nu,i}^{z}\right\rangle=0$, but $\left\langle S_{1,i}^{z}S_{2,i}^{z}\right\rangle\neq 0$. The transition between the ferromagnet and the product phase is also of infinite-randomness type surrounded by quantum Griffiths ferromagnetic and product phases. By duality, [^1] The Hamiltonian (1) is invariant under the duality transformation: $S_{\nu,i}^{z}S_{\nu,i+1}^{z}\rightarrow\tau_{\nu,i}^{x}$, $S_{\nu,i}^{x}\rightarrow\tau_{\nu,i}^{z}\tau_{\nu,i+1}^{z}$ (here, $\tau^{x}$ and $\tau^{z}$ are the dual Pauli operators), $J_{i}\leftrightarrow h_{i}$, and $K_{i}\leftrightarrow g_{i}$. the transition between the paramagnetic and product phases is in the same universality class (2014_ashkinteller).
@@
From the first and second sums of the Hamiltonian (4), it is clear that, when $\epsilon_{I}\gg 1$ and sufficiently near to the self-duality line $\left(\delta\approx 0\right)$, the system is in a phase with long-range order in the $\boldsymbol{\sigma}$ variables ($\left\langle\sigma^{z}\right\rangle\neq 0$) and disordered in the $\boldsymbol{\eta}$ variables ($\left\langle\eta^{z}\right\rangle=0$). This implies that long-range order develops in the product variables $\left\langle S_{1}^{z}S_{2}^{z}\right\rangle\neq 0$, while the spins remain disordered in each individual color $\left\langle S_{1}^{z}\right\rangle=\left\langle S_{2}^{z}\right\rangle=0$ (1943_ashkinteller; 1981_ashkinteller). Thus, the moniker product (PROD) phase. A caveat is needed here. The terminology $\boldsymbol{\sigma}_{i}$ and $\boldsymbol{\eta}_{i}$ as a product and original variables is due to the lack of a better one. It is precise if one considers only the $z$-component of the original spins [see Eq. (2)]. Regarding the $x$-component [see Eq. (3)], however, the terminology is the other way round. Regarding the $y$-component, both could be regarded as product variables. In sum, both $\boldsymbol{\sigma}_{i}$ and $\boldsymbol{\eta}_{i}$ are nontrivial compositions of the original spin variables. We stick with this terminology because the $z$-component is the one defining the order parameter field of the transition.
@@
In the presence of quenched disorder, it was found that all quantum phase transitions belong to the same universality class (that of the random transverse-field Ising chain) and are surrounded by the associated quantum Griffiths phases (2014_ashkinteller) as illustrated in the phase diagram Fig. 1(a).
@@
Figure 1: (a) Schematic phase diagram of the RQAT chain without dissipation. The dashed lines indicate crossovers from conventional FM, PM and product phases to Griffiths phases. GFM and GPM denote the Griffiths ferromagnetic and paramagnetic phases, respectively. DG corresponds to the double Griffiths phase, characterized by the presence of rare regions in both $\sigma$ and $\eta$ variables. GPROD denotes the Griffiths product phase, where rare regions form in the $\sigma$ variable near the PM transition and in the $\eta$ variable near the FM transition. (b) Schematic phase diagram of the RQAT chain with dissipation. The transitions where rare regions are formed in the original spin variable are smeared. In the strong-coupling regime, the GFM phase is replaced by a weakly ordered (WO) phase and the GPROD phase is replaced by an inhomogeneous ferromagnetic (IFM) phase.
@@
It is important to distinguish between the two sorts of Griffiths phases in the region $\epsilon_{I}>1$. Surrounding the PM–PROD transition, the Griffiths singularities in both the Griffiths PM and Griffiths PROD phases are due to rare regions locally in a PROD phase and weakly interacting with the bulk. In RRs of this kind, the $z$-component of the original spins fluctuate rapidly and independently of the nearby spins. However, they fluctuate together with the other color spin in the same site. On the other hand, surrounding the PROD-FM transition, the low-energy behavior of both the Griffiths PROD and FM phases are dominated by RRs locally in the FM phase and weakly coupled to the bulk. Near the multicritical point, both RRs coexist in the region $\epsilon_{I}>1$ and, therefore, give rise to the double Griffiths (DG) phase (2014_ashkinteller). It is worth noting that this DG phase have been seen in a density-matrix renormalization-group study (chatelain-voliotis-epjb16).
@@
Evidently, these two phase transitions and the associated Griffiths phases are related by duality. However, this will be broken when dissipation is introduced as it couples to the original spin variables and not to the product ones. As a result, only the FM-like RRs will be dramatically affected by dissipation yielding to the smearing of the transition as sketched in Fig. c.
III A simple dissipative model
@@
where $H_{\text{S}}$ is the RQAT Hamiltonian defined in Eq. (1). The bath component are independent harmonic oscillators, and thus,
@@
Here, the oscillator frequencies $\omega_{k,i}$ and masses $m_{k,i}$ are assumed to be color-independent. Finally, following the Caldeira-Leggett model (caldeiraleggett; leggett_spinboson; smeared_hoyos_1; smeared_hoyos_2), the system-bath coupling is
@@
where the constant $\lambda_{k,i}$ controls the strength of the coupling. For the moment, we consider color-independent coupling constants. Later, we discuss the effects of color-dependence.
@@
In what follows, we follow Refs. (smeared_hoyos_1; smeared_hoyos_2) to generalize the adiabatic renormalization procedure (leggett_spinboson) to the present case. Importantly, we assume weak dissipation ($\alpha_{i}\ll 1$). This means that, individually, each spin is weakly affected by dissipation. Moreover, we also assume that the phases of the undamped system are not unstable against weak dissipation, i.e., we assume that the nature of the phases do not change by weak dissipation alone.
@@ IV Adiabatic renormalization
This section applies the adiabatic renormalization procedure (leggett_spinboson) to our model system. We remark that it is applicable to the regimes of Ohmic and super-Ohmic dissipation $\left(s\geq 1\right)$. Although not applicable to the sub-Ohmic regime ($s<1$), insights can be gained.
@@
where an irrelevant constant term has been omitted for simplicity and $a_{0;k,i}=\lambda_{k,i}/m_{k,i}\omega_{k,i}^{2}$ is a constant. Notice that $H_{0}$ is a collection independent harmonic oscillators with offset rest positions depending on $\eta_{i}^{z}$ and the other on $\sigma_{i}^{z}\eta_{i}^{z}$ and, thus, its ground state is fourfold degenerate:
$\begin{array}[]{cc}\left|\Psi_{\pm\pm}\right\rangle=&\left|\pm,\pm\right\rangle\bigotimes_{\omega_{k,i}>\omega_{l}}\left|G_{k,i;\pm}^{(1)}\right\rangle\left|G_{k,i;\pm\pm}^{(2)}\right\rangle,\\ \left|\Psi_{\pm\mp}\right\rangle=&\left|\pm,\mp\right\rangle\bigotimes_{\omega_{k,i}>\omega_{l}}\left|G_{k,i;\pm}^{(1)}\right\rangle\left|G_{k,i;\pm\mp}^{(2)}\right\rangle,\end{array}$ (11)
where we are using the notation $\left|\eta_{i}^{z},\sigma_{i}^{z}\right\rangle$. The oscillators ground states have the form
@@
with $\left|0_{k,i}^{(\nu)}\right\rangle$ denoting the ground state of a oscillator with rest position at the origin. The tunneling term in the Hamiltonian (4) is
@@
The other terms are all diagonal in the ground state and can simply be incorporated to $H_{0}$ latter on. Treating $H_{1}$ perturbatively up to first order, we get the effective Hamiltonian
@@
$\tilde{h}_{i}=h_{i}e^{-\frac{2}{\pi}\intop_{p\epsilon_{h,i}h_{i}}^{\omega_{c}}d\omega\frac{F_{i}(\omega)}{\omega^{2}}}=h_{i}e^{-\alpha_{i}\intop_{p\epsilon_{h,i}h_{i}}^{\omega_{c}}d\omega\frac{\omega^{s-2}}{\omega_{c}^{s-1}}},$ (16)
@@
$\tilde{g}_{i}=g_{i}e^{-\frac{4}{\pi}\intop_{pg_{i}}^{\omega_{c}}d\omega\frac{F_{i}(\omega)}{\omega^{2}}}=g_{i}e^{-2\alpha_{i}\intop_{pg_{i}}^{\omega_{c}}d\omega\frac{\omega^{s-2}}{\omega_{c}^{s-1}}}.$ (17)
@@
Because the effective Hamiltonian (15) keeps its original form, this procedure can be iterated until convergence $g_{i}^{*}$ and $h_{i}^{*}$. As shown in Ref. (leggett_spinboson), in the Ohmic regime ($s=1$), $h_{i}^{*}$ and $g_{i}^{*}$ converge to finite values $h_{i}(p\epsilon_{h}h_{i}/\omega_{c})^{\alpha/\left(1-\alpha\right)}$ and $g_{i}(pg_{i}/\omega_{c})^{2\alpha/\left(1-2\alpha\right)}$ if $\alpha_{i}<1$ and <1/2<1/2, respectively, otherwise, they converge to vanishing values, and tunneling ceases. For super-Ohmic dissipation ($s>1$), $g_{i}^{*}$ and $h_{i}^{*}$ are always finite.
As we are interested in the case of weak dissipation, the bare value of the dissipation parameter $\alpha_{i}\ll 1$. Therefore, dissipation has little effect individually as $h_{i}^{*}\approx h_{i}$ and $g_{i}^{*}\approx g_{i}$. However, near a phase transition, spins at different sites become correlated and, as shown in Ref. (smeared_hoyos_1), the effective dissipation parameter can change under RG. The central question then becomes the following: under what conditions does the dissipation parameter flow to larger values? As we show in the following, this happens only when FM-like RRs appear, and in the FM phase. Surprisingly, not when PROD-like RRs are formed.

### v2

@@ I Introduction
Quenched disorder underlies a manifold of profound effects near quantum phase transitions: For instance, disorder-induced inhomogeneities lead to exotic universality classes such as those typified by the infinite randomness critical point [1, 2], and the strong off-critical Griffiths singularities [3, 4]. Furthermore, in certain situations these inhomogeneities could lead to the smearing [5, 6] of the sharp quantum phase transition. The moorings of the effects described above is centered around the so-called rare regions (RRs): arbitrarily large regions in space which, due to a statistically rare configuration of the quenched-disordered couplings, is locally ordered and weakly coupled to the bulk. These RRs, characterized by slow dynamics, dominate the low-energy behavior of physical observables [7, 8], (for reviews, see Refs. [9, 10, 11, 12]).
@@
In practice it is important to realize that real materials are seldom described as closed systems. Thus it is of paramount importance to understand how the phases and phase transitions get affected when the system is connected to a dissipative reservoir. Initially, it was realized that sufficiently strong dissipation can completely damp the fluctuations of a RR [13, 14, 15, 16], thus establishing that the combined effect of dissipation and disorder is nonperturbative and could round the sharp phase transition by smearing. To quantify the resulting phase and phase transition, interactions among the RRs should be taken into account and it was shown that this could be accomplished via a the strong-disorder renormalization-group (SDRG) method [17, 18]. Later, it was established that for the case of the random transverse-field Ising model (RTFIM), in any dimension, the critical point is entirely destroyed by smearing when dissipation is Ohmic or Sub-Ohmic [5, 6]. Physically, the smearing phenomenon occurs because dissipation freezes the dynamics of sufficiently large RRs, independent of the bulk, i.e., each rare region orders independently from the bulk, leading to local phase transitions that can take place at different values of the control parameter. Thus the paramagnetic Griffiths phase becomes an inhomogeneously ordered ferromagnetic phase. Many other numerical studies (either SDRG or Monte Carlo) have confirmed this conclusion [11, 12]. Notably, experimental evidence of smeared phase transitions has been observed in the compounds Sr1-xCaxRuO3 [19] and CePd1-xRhx [20]. More recently, smeared phase transitions have also been identified in a numerical study of strange metals [21], and have also potentially been seen in low-dimensional superconductors [22].
@@
This work belongs to the genre of problems that we have discussed in the previous paragraph: Namely, we study the dissipative random quantum Ashkin-Teller (RQAT) model. Besides the usual paramagnetic (PM) and ferromagnetic (FM) phases, this model features an additional phase, called the product phase, which exhibits a composite (or intertwined) order [23]. We inquire whether the transitions between this product phase and the other phases are smeared by the combined effects of dissipation and disorder. Interestingly, we show that Ohmic dissipation smears only the transition between the product and FM phases, while the transition between product and PM phases remains sharp. This conclusion is robust and persists across different forms of system-bath couplings. Surprisingly, even though dissipation is coupled to the original order parameter fields, it does not couple generically to the intertwined order parameter field and, therefore, does not damp the quantum fluctuations of the associated rare regions that form near the transition line between the product and PM phases.
@@
Finally, besides this fundamental motivation, it is worth mentioning that the Ashkin-Teller model has applications in other contexts such as to describe layers of atoms absorbed on surfaces [24], current loops in high-$T_{c}$ superconductors [25], and the elastic response of DNA molecules [26]. The understanding of dissipation in these contexts is another motivation for our work.
@@
In the remainder of this work, we review the effects of disorder on the quantum Ashkin-Teller chain in Sec. II, introduce the model studied in Sec. III, and apply the adiabatic renormalization-group and the strong-disorder renormalization-group methods to our problem in Secs. IV and V, respectively. The resulting phase diagram is then discussed in Sec. VI. Finally, we give concluding remarks in Sec. VII. Other possible models of dissipation are considered in the Appendix A and we find that our conclusions remain qualitatively correct.
II Review of the dissipationless model
II.1 Original variables
@@
The Hamiltonian of the one-dimensional random quantum Ashkin-Teller (RQAT) model is [27, 23, 28, 29, 30]
@@
where $S_{\nu,i}^{z}$ and $S_{\nu,i}^{x}$ are Pauli matrices representing local spin-1/2 moments. The first two terms on the right-hand-side of Eq. (1) describe two independent random transverse-field Ising (RTFI) chains whereas the remaining terms quantify the couplings between them. The index $\nu=1,2$, referred to as the color index, distinguishes the two chains and the index $i$ distinguishes different lattice sites. The interactions $J_{i}$ and transverse fields $h_{i}$ are strictly positive, independent random variables. The coupling strengths are characterized by the ratios $\epsilon_{h,i}=g_{i}/h_{i}$ and $\epsilon_{J,i}=K_{i}/J_{i}$. For simplicity, we assume that the (bare) values $\epsilon_{h,i}=\epsilon_{J,i}=\epsilon_{I}>0$; i.e., they are uniform and site independent. However, they flow under the renormalization-group transformations and become random variables [30]. The low-energy physical behavior of the model (1) can be divided into two regimes: the weak-coupling regime ($\epsilon_{I}<1$) and the strong-coupling regime ($\epsilon_{I}>1$). In the low-energy limit of the weak-coupling regime, the two Ising chains effectively decouple [28, 29] and the resulting physics is that of two identical RTFI chains. Therefore, there is a direct quantum phase transition between a ferromagnet (where the quantum average $\left\langle S_{\nu,i}^{z}\right\rangle\neq 0$) and paramagnet (where $\left\langle S_{\nu,i}^{z}\right\rangle=0$). The transition, which occurs when $\prod J_{i}=\prod h_{i}$, is of infinite-randomness type and is surrounded by the Griffiths ferro- and paramagnetic phases. In the strong-coupling regime $\epsilon_{I}>1$, on the other hand, this transition is through an intervening product phase exhibiting a composite order; i.e., $\left\langle S_{\nu,i}^{z}\right\rangle=0$, but $\left\langle S_{1,i}^{z}S_{2,i}^{z}\right\rangle\neq 0$. The transition between the ferromagnet and the product phase is also of infinite-randomness type surrounded by quantum Griffiths ferromagnetic and product phases. By duality, [^1] The Hamiltonian (1) is invariant under the duality transformation: $S_{\nu,i}^{z}S_{\nu,i+1}^{z}\rightarrow\tau_{\nu,i}^{x}$, $S_{\nu,i}^{x}\rightarrow\tau_{\nu,i}^{z}\tau_{\nu,i+1}^{z}$ (here, $\tau^{x}$ and $\tau^{z}$ are the dual Pauli operators), $J_{i}\leftrightarrow h_{i}$, and $K_{i}\leftrightarrow g_{i}$. the transition between the paramagnetic and product phases is in the same universality class [30].
II.2 Product variable
@@
From the first and second sums of the Hamiltonian (4), it is clear that, when $\epsilon_{I}\gg 1$ and sufficiently near to the self-duality line $\left(\delta\equiv\left\langle\ln\left(h/J\right)\right\rangle\approx 0\right)$, the system is in a phase with long-range order in the $\boldsymbol{\sigma}$ variables ($\left\langle\sigma^{z}\right\rangle\neq 0$) and disordered in the $\boldsymbol{\eta}$ variables ($\left\langle\eta^{z}\right\rangle=0$). This implies that long-range order develops in the product variables $\left\langle S_{1}^{z}S_{2}^{z}\right\rangle\neq 0$, while the spins remain disordered in each individual color $\left\langle S_{1}^{z}\right\rangle=\left\langle S_{2}^{z}\right\rangle=0$ [27, 23]- thus, the moniker “product (PROD) phase”. A caveat is needed here. The terminology $\boldsymbol{\sigma}_{i}$ and $\boldsymbol{\eta}_{i}$ as a product and original variables is due to the lack of a better one. It is precise if one considers only the $z$ component of the original spins [see Eq. (2)]. Regarding the $x$ component [see Eq. (3)], however, the terminology is the other way round. Regarding the $y$ component, both could be regarded as product variables. In sum, both $\boldsymbol{\sigma}_{i}$ and $\boldsymbol{\eta}_{i}$ are nontrivial compositions of the original spin variables. We stick with this terminology because the $z$ component is the one defining the order parameter field of the transition.
II.3 Phase diagram: Dissipationless case
@@
In the presence of quenched disorder, it was found that all quantum phase transitions belong to the same universality class (that of the random transverse-field Ising chain) and are surrounded by the associated quantum Griffiths phases [30] as illustrated in the phase diagram Fig. 1(a).
@@
Figure 1: (a) Schematic phase diagram of the RQAT chain without dissipation ($\alpha=0$, discussed in Sec. II). GFM and GPM denote the Griffiths ferromagnetic and paramagnetic phases, respectively. DG corresponds to the double Griffiths phase, characterized by the presence of rare regions in both $\sigma$ and $\eta$ variables. GPROD denotes the Griffiths product phase, where rare regions form in the $\sigma$ variable near the PM transition and in the $\eta$ variable near the FM transition. (b) Schematic phase diagram of the RQAT chain with dissipation ($\alpha\neq 0$, discussed in Sec. VI). The transitions where rare regions are formed in the original spin variable are smeared. In the strong-coupling regime, the GFM phase is replaced by a weakly ordered (WO) phase and the GPROD phase is replaced by an inhomogeneous ferromagnetic (IFM) phase. In all cases, thick solid lines represent second-order quantum phase transitions. The dot-dashed line in (b) represents the smeared transition between the PROD and FM phases. The dashed lines indicate crossovers, not transitions. Different colors represent different phases.
@@
It is important to distinguish between the two sorts of Griffiths phases in the region $\epsilon_{I}>1$. Surrounding the PM-PROD transition, the Griffiths singularities in both the Griffiths PM and Griffiths PROD phases are due to rare regions locally in a PROD phase and weakly interacting with the bulk. In RRs of this kind, the $z$ components of the original spins fluctuate rapidly and independently of the nearby spins. However, they fluctuate together with the other color spin in the same site. On the other hand, surrounding the PROD-FM transition, the low-energy behavior of both the Griffiths PROD and FM phases is dominated by RRs locally in the FM phase and weakly coupled to the bulk. Near the multicritical point, both RRs coexist in the region $\epsilon_{I}>1$ and, therefore, give rise to the double Griffiths (DG) phase [30]. It is worth noting that this DG phase has been seen in a density-matrix renormalization-group study [31].
@@
Evidently, these two phase transitions and the associated Griffiths phases are related by duality. However, this will be broken when dissipation is introduced as it couples to the original spin variables and not to the product ones. As a result, only the FM-like RRs will be dramatically affected by dissipation yielding to the smearing of the transition as sketched in Fig. 1(b) (see further discussions in Sec. VI).
III Simple dissipative model
@@
where $H_{\text{S}}$ is the RQAT Hamiltonian defined in Eq. (1). The bath components are independent harmonic oscillators and, thus,
@@
Here, the oscillator frequencies $\omega_{k,i}$ and masses $m_{k,i}$ are assumed to be color independent. Finally, following the Caldeira-Leggett model [32, 33, 5, 6], the system-bath coupling is
@@
where the constant $\lambda_{k,i}$ controls the strength of the coupling. For the moment, we consider color independent coupling constants. Later, we discuss the effects of color dependence.
@@
In what follows, we follow Refs. [5, 6] to generalize the adiabatic renormalization procedure [33] to the present case. Importantly, we assume weak dissipation ($\alpha_{i}\ll 1$). This means that, individually, each spin is weakly affected by dissipation. Moreover, we also assume that the phases of the undamped system are not unstable against weak dissipation, i.e., we assume that the nature of the phases does not change by weak dissipation alone.
@@ IV Adiabatic renormalization
This section applies the adiabatic renormalization procedure [33] to our model system. We remark that it is applicable to the regimes of Ohmic and super-Ohmic dissipation $\left(s\geq 1\right)$. Although not applicable to the sub-Ohmic regime ($s<1$), insights can be gained.
@@
where an irrelevant constant term has been omitted for simplicity and $a_{0;k,i}=\lambda_{k,i}/m_{k,i}\omega_{k,i}^{2}$ is a constant. Notice that $H_{0}$ is a collection of independent harmonic oscillators with offset rest positions depending on $\eta_{i}^{z}$ and the other on $\sigma_{i}^{z}\eta_{i}^{z}$ and, thus, its ground state is fourfold degenerate:
$\left|\Psi_{\pm\pm}\right\rangle$ $=$ $\left|\pm,\pm\right\rangle\bigotimes_{\omega_{k,i}>\omega_{l}}\left|G_{k,i;\pm}^{(1)}\right\rangle\left|G_{k,i;\pm\pm}^{(2)}\right\rangle,$ (11)
$\left|\Psi_{\pm\mp}\right\rangle$ $=$ $\left|\pm,\mp\right\rangle\bigotimes_{\omega_{k,i}>\omega_{l}}\left|G_{k,i;\pm}^{(1)}\right\rangle\left|G_{k,i;\pm\mp}^{(2)}\right\rangle,$ (12)
where we are using the notation $\left|\eta_{i}^{z},\sigma_{i}^{z}\right\rangle$. The oscillators’ ground states have the form
@@
with $\left|0_{k,i}^{(\nu)}\right\rangle$ denoting the ground state of an oscillator with rest position at the origin. The tunneling term in the Hamiltonian (4) is
@@
The other terms are all diagonal in the ground state and can simply be incorporated to $H_{0}$ later on. Treating $H_{1}$ perturbatively up to first order, we get the effective Hamiltonian
@@
$\tilde{h}_{i}=h_{i}e^{-\frac{2}{\pi}\intop_{p\epsilon_{h,i}h_{i}}^{\omega_{c}}d\omega\frac{F_{i}(\omega)}{\omega^{2}}}\approx h_{i}e^{-\alpha_{i}\intop_{p\epsilon_{h,i}h_{i}}^{\omega_{c}}d\omega\frac{\omega^{s-2}}{\omega_{c}^{s-1}}},$ (17)
@@
$\tilde{g}_{i}=g_{i}e^{-\frac{4}{\pi}\intop_{pg_{i}}^{\omega_{c}}d\omega\frac{F_{i}(\omega)}{\omega^{2}}}\approx g_{i}e^{-2\alpha_{i}\intop_{pg_{i}}^{\omega_{c}}d\omega\frac{\omega^{s-2}}{\omega_{c}^{s-1}}}.$ (18)
@@
Because the effective Hamiltonian (16) keeps its original form, this procedure can be iterated until convergence of $g_{i}^{*}$ and $h_{i}^{*}$. As shown in Ref. [33], in the Ohmic regime ($s=1$), $h_{i}^{*}$ and $g_{i}^{*}$ converge to finite values $h_{i}(p\epsilon_{h}h_{i}/\omega_{c})^{\alpha/\left(1-\alpha\right)}$ and $g_{i}(pg_{i}/\omega_{c})^{2\alpha/\left(1-2\alpha\right)}$ if $\alpha_{i}<1$ and <1/2<1/2, respectively, otherwise, they converge to vanishing values, and tunneling ceases. For super-Ohmic dissipation ($s>1$), $g_{i}^{*}$ and $h_{i}^{*}$ are always finite.
As we are interested in the case of weak dissipation, the bare value of the dissipation parameter $\alpha_{i}\ll 1$. Therefore, dissipation has little effect individually as $h_{i}^{*}\approx h_{i}$ and $g_{i}^{*}\approx g_{i}$. However, near a phase transition, spins at different sites become correlated and, as shown in Ref. [5], the effective dissipation parameter can change under RG. The central question then becomes the following: under what conditions does the dissipation parameter flow to larger values? As we show in the following, this happens only when FM-like RRs appear, and in the FM phase - surprisingly, not when PROD-like RRs are formed.